OpenStudy (butterflyhope):

Urgent *will medal*: If y = 3x - 5 and x = 3y - 5, then what are x and y? Thanks c:

OpenStudy (bohotness):

i go t it :)

OpenStudy (anonymous):

y = 3x - 5 x = 3y - 5 Put y = 3x-5 in 2nd eq x = 3(3x-5) -5 Find x then use that value to find y

OpenStudy (bohotness):

x=3y-5 y=3x-5